What is Cramer's V for each of the following values for the chi-square test for independence? (Round your answers to two decimal places.)

(a)    X2 = 8.44, n = 130, dfsmaller = 2

V =  

(b)    X2 = 5.17, n = 50, dfsmaller = 1

V =  

(c)    X2 = 11.76, n = 190, dfsmaller = 3

V =
Answer #1

a)

V= sqrt ( 8.44 / 130*2)

= 0.18

b)

V= sqrt ( 5.17/ 50*1)

= 0.32

c)

V= sqrt ( 11.76/ 190*3)

=0.14


Please revert back in case of any doubt.

Please upvote. Thanks in advance.